OpenStudy (anonymous):

Kara read 30 pages of a book in 37 minutes. how long will it take her to read the remaining 247 pages in the book, in minutes?

OpenStudy (anonymous):

30 pages = 37 minutes 1 page = 37/30 minutes 247 pages = x minutes \[247 pages = 247 pages \frac{37 minutes}{30 pages}=\frac{247*37}{30}=304.63333=305 \,minutes\]